# Uni3C: Unifying Precisely 3D-Enhanced Camera and Human Motion Controls for Video Generation

### v1

@@ 5.4 Ablation Study and Exploratory Discussions
The point clouds of humans are always frozen in the world space without any motion, which creates a conflict with the human motion conditions provided by SMPL-X. Eliminating this “redundant” information is a straightforward idea to improve motion quality However, as verified in Table 2, while Uni3C-T2V with human-masked point clouds achieves slightly better visual quality, this masking adversely affects camera precision, particularly when humans occupy a significant area of the image. Therefore, retaining the point clouds of humans is essential for effective camera control. Given that Wan2.1 [54] is trained on videos featuring substantial motion, it can generate natural and smooth movements even when the foreground point clouds remain fixed.
Figure 7: Rendering results with and without gravity calibration by GeoCalib [53].
@@
We present results of motion transfer achieved by the Uni3C framework in Figure 13. Our model effectively controls both camera trajectories and human motions, even when reference motions are sourced from different videos or distinct domains, such as animation and real-world scenes. Meanwhile, Uni3C can be further extended to generate vivid videos based on other conditions, like text-to-motion guidance or retrieved motions from motion databases. To prove this point, we randomly integrate several motion clips from BABEL [41] and use Uni3C to control both motion and camera, as illustrated in Figure 14. More results are shown on our projected page.
Limitation.
@@
Although Uni3C supports flexible and diverse unified control and motion transfer, it operates under the constraints of predefined camera trajectories and human characters (SMPL-X). Consequently, Uni3C may struggle to produce physically plausible outcomes when human motions conflict with environmental conditions, as illustrated in Figure 15. For instance, if a human’s movement trajectory is blocked by walls, barriers, or other objects, the generated results may exhibit artifacts such as distortion, clipping, or sudden disappearance. This limitation could be mitigated by employing a more advanced human motion generation method that accounts for physical obstructions within the environment.
6 Conclusion
@@
This paper introduced Uni3C, a framework that unifies 3D-enhanced camera and human motion controls for video generation. We first propose the PCDController, demonstrating that lightweight, trainable modules, and rich geometric priors from 3D point clouds can efficiently manage camera trajectories without compromising the inherent capacities of foundational VDMs. This not only enhances generalization but also facilitates versatile downstream applications without joint training. Furthermore, by aligning multi-modal conditions, including both environmental point clouds and human characters in a global 3D world space, we established a coherent framework for jointly controlling camera movements and human animations. Our comprehensive experiments validated the efficacy of Uni3C across diverse datasets, showcasing superior performance in both quantitative and qualitative assessments compared to existing approaches. The significance of our contributions lies not only in improving the state-of-the-art in controllable video generation but also in proposing a robust way to inject multi-modal conditions without the requirements of heavily annotated data. We believe that Uni3C paves the way for more advanced controllable video generation.
Figure 8: Our out-of-distribution benchmark for camera control. The validation set includes generative, human, scene-level, and object-level images with diverse aspect ratios.
@@
Figure 9: Qualitative results of camera control on our benchmark. We compare the proposed PCDController to ViewCrafter [66], SEVA [71], and our model without point cloud guidance. The leftmost image is the reference condition. “full” indicates using both Plücker ray and point clouds as conditions.
@@
Figure 10: Results of the challenging orbital 360 rotations from PCDController. The leftmost images are the reference views.
@@
Figure 11: Results of unified camera and human motion controls. The leftmost images are the reference views, while the first row indicates the aligned 3D world guidance.
@@
Figure 12: Results of unified camera, human motion, and Hamer controls. The leftmost images are the reference views, while the first and second rows indicate the aligned 3D world guidance and Hamer rendering, respectively.
@@
Figure 13: Results of motion transfer. The first row indicates the reference video, while others show our generated videos transferring motions from the reference sequence.
@@
Figure 14: Results transferred from random integrated motion clips of BABEL [41]. The motion sequences are listed on the left, which are executed from light to dark colors.
@@
Figure 15: Failed cases generated by Uni3C. These results are primarily limited by the conflict between human motions and environments.

### v2

@@ 5.4. Ablation Study and Exploratory Discussions
Table 4. Ablation of PCDController. Different DiT layers are employed from top to bottom. PCDController adopts 20-layer as the final setting.
[…]
Layer Param Score
10 0.48B 87.75 0.136 0.0398 0.559
20 0.95B 88.27 0.102 0.0313 0.246
30 1.39B 88.08 0.092 0.0302 0.197
40 1.85B 87.90 0.115 0.0339 0.216
@@
Human point clouds are always frozen in world space without any motion, which conflicts with the human motion conditions provided by SMPL-X. Eliminating this “redundant” information is a straightforward idea to improve motion quality. However, as verified in Table 3, while Uni3C-T2V with human-masked point clouds achieves slightly better visual quality, this masking adversely affects camera precision, particularly when humans occupy a significant image area. Therefore, retaining the point clouds of humans is essential for effective camera control. Given that Wan2.1 (Wang et al., 2025a) is trained on videos featuring substantial motion, it can generate natural and smooth movements even when the foreground point clouds remain fixed.
@@
We present motion transfer results achieved by the Uni3C framework in Figure 11. Our model effectively controls both camera trajectories and human motions, even when reference motions are sourced from different videos or distinct domains, such as animation and real-world scenes. Meanwhile, Uni3C can be further extended to generate vivid videos based on other conditions, like text-to-motion guidance or retrieved motions from motion databases. To prove this point, we randomly integrate several motion clips generated text-to-motion (Barquero et al., 2024) trained on BABEL (Punnakkal et al., 2021) and use Uni3C to control both motion and camera, as illustrated in Figure 12. More results are shown in our supplementary.
Ablation of PCDController.
We present some ablation results of PCDController with various external DiT layers in Table 4. Each setting injects conditional features into the main Wan-I2V backbone from top to bottom with different layers. While the DiT branch with 30 layers shows slightly improved camera metrics, it ultimately compromises overall visual quality and generalization. We reveal that solely applying more external layers suffers from overfitting point cloud conditions as shown in Figure 15, resulting in suboptimal outcomes with distorted point clouds. Thus, we determine that employing 20 DiT layers achieves an effective balance between controllability and generalization within the PCDController.
5.5. User Study
We conducted user studies for camera control (128 samples) and unified control (150 samples) on our respective benchmarks in Table 5. Formally, 20 unrelated volunteers are invited. For the camera control, volunteers are requested to vote for the best result, considering video quality and camera precision, while the quality of human pose is additionally required for the unified control. If two methods perform very similarly, volunteers are allowed to vote for both of them as top-2. We finally average the score for each metric.
Table 5. User studies on camera and unified (camera&human) control.
Camera CamI2V ViewCrafter SEVA PCDController
[…]
Camera&Human CamAnimate RealisDance-DiT PCDController Uni3C
[…]
6. Limitation and Future Work
@@
Although Uni3C supports flexible and diverse unified control and motion transfer, it operates under the constraints of predefined camera trajectories and human characters (SMPL-X). Consequently, Uni3C struggles to produce physically plausible outcomes when human motions conflict with environmental conditions, as shown in Figure 16. For instance, if a human’s movement trajectory is blocked by walls, barriers, or other objects, the generated results may exhibit artifacts such as distortion, clipping, or sudden disappearance. This limitation could be mitigated by employing a more advanced human motion generation method that accounts for physical obstructions within the environment. Besides, though our method can produce plausible results within inaccurate SMPL or world coordinates, it will result in undesired pose or motion direction.
7. Conclusion
@@
This paper introduced Uni3C, a framework that unifies 3D-enhanced camera and human motion controls for video generation. We first propose the PCDController, demonstrating that lightweight, trainable modules, and rich geometric priors from 3D point clouds can efficiently manage camera trajectories without compromising the inherent capacities of foundational VDMs. This not only enhances generalization but also facilitates versatile downstream applications without joint training. Furthermore, by aligning multi-modal conditions, including both environmental point clouds and human characters in a global 3D world space, we established a coherent framework for jointly controlling camera movements and human animations. Our comprehensive experiments validated the efficacy of Uni3C across diverse datasets, showcasing superior performance in both quantitative and qualitative assessments compared to existing approaches. The significance of our contributions lies not only in improving the state-of-the-art in controllable video generation but also in proposing a robust way to inject multi-modal conditions without the requirements of heavily annotated data. We believe that Uni3C paves the way for advanced controllable video generation.
Acknowledgements.
This work was supported by the Science and Technology Commission of Shanghai Municipality (No. 24511103100).
@@ Acknowledgements.
Figure 7. Qualitative results of camera control on our benchmark. We compare the proposed PCDController to ViewCrafter (Yu et al., 2024), SEVA (Zhou et al., 2025a), and our model without point cloud guidance. The leftmost image is the reference condition. “full” indicates using both Plücker ray and point clouds as conditions.
@@
Figure 8. Results of the challenging orbital 360 rotations from PCDController. The leftmost images are the reference views.
@@
Figure 9. More results of unified camera and human motion controls. Leftmost images are reference views; the first row indicates aligned 3D world guidance.
@@
Figure 10. Results of unified camera, human motion, and Hamer controls. The leftmost images are the reference views, while the first and second rows indicate the aligned 3D world guidance and Hamer rendering.
@@
Figure 11. Results of motion transfer. The first row indicates the reference video, while others show our generated videos transferring motions from the reference sequence.
@@
Figure 12. Results transferred from randomly integrated motion clips generated from text-to-motion (Barquero et al., 2024) trained on BABEL (Punnakkal et al., 2021). The motion sequences are listed on the left, which are executed from light to dark colors.
Figure 13. Ablation results of Plücker ray and point clouds during training phase. Point clouds enjoy highly accurate camera control against Plücker ray.
Figure 14. Rendering results with and without gravity calibration by GeoCalib (Veicht et al., 2024).
Figure 15. Camera control results of PCDController with challenging point clouds. The model with fewer external DiT layers enjoys superior conditional robustness compared to the one with more DiT layers.
@@
Figure 16. Failed cases generated by Uni3C. These results are primarily limited by the conflict between human motions and environments.
Figure 17. Our out-of-distribution benchmark for camera control. The validation set includes generative, human, scene-level, and object-level images with diverse aspect ratios.
Appendix A Border Impact
This study delves into the realm of controllable video generation. The remarkable generative capabilities of AI-generated content (AIGC) can inadvertently lead to the creation of misleading information or fabricated visuals. Consequently, we sincerely urge users to remain vigilant regarding these issues. Moreover, issues of privacy and consent must be taken into account, as generative models are always developed using extensive datasets. It’s also crucial to acknowledge that such models can potentially reinforce biases in the training data, which may result in unjust outcomes. Therefore, we advise users to be responsible and inclusive when utilizing these generative models. It is important to mention that our method concentrates solely on technical aspects. All pre-trained models and training videos referenced in this study are publicly available.
Appendix B Details of Uni3C
Figure 18. Detailed model architecture of PCDController.
B.1. Architecture Details of PCDController
As mentioned in the main paper, our PCDController is built upon a simplified DiT module. We apply detailed ablation studies and demonstrate that the 20-layer external DiT branch with 1024 hidden size enjoys a good balance between both controllability and generalization. Detailed architecture of PCDController is illustrated in Figure 18. We follow the DiT design in CogVideoX (Yang et al., 2025), including AdaLayerNorm, 3D attention, and Feed Forward Network (FFN), without the textual branch. The output features are added to the main branch of Wan-I2V block by block via zero-initialized projection layers.
Figure 19. Detailed architecture of Realisdance-DiT (Zhou et al., 2025b).
B.2. Architecture Details of Realisdance-DiT
We show the model architecture of Realisdance-DiT (Zhou et al., 2025b) in Figure 19 employed for our unified controls for both camera and human poses. Formally, Realisdance-DiT is fine-tuned from the pre-trained Wan-I2V, while features of human condition signals (SMPL-X and Hamer) are added to the inputs, and the reference image feature is spatially concatenated to the inputs to confirm fine-grained details. During the training of Realisdance-DiT, only convolution-based pose and reference patchfiers and self-attention layers of DiT are trainable.
B.3. High-Resolution Inference
While PCDController is trained under 480p multi-view images ($480\times 768$, $512\times 720$, $608\times 608$, $720\times 512$, $768\times 480$), we find that this model also performs well under 720p video generation ($720\times 1280$, $800\times 1152$, $960\times 960$, $1152\times 800$, $1280\times 720$) without specific fine-tuning, as empirically verified in our qualitative results in the supplementary. Specifically, for the inference of 720p, we first render the conditional videos of warped point clouds with PyTorch3D under 480p. Then we replace the first frame with a 720p reference image as the high-resolution guidance. Thus, PCDController can produce high-quality 720p videos extended from the high-resolution reference, while the low-resolution point cloud renderings are only considered as auxiliary camera signals.
B.4. Inference Efficiency
Table 6. Inference efficiency of Uni3C. Our model only increases a little inference time compared to the baseline method.
Methods Inference Times
[…]
RealisDance-DiT 50.5s 182.7s
Uni3C 62.3s (+23%) 213.4s (+17%)
Table 7. Efficiency of other components of Uni3C.
PyTorch3D-Rendering 2D-Keypoint+Alignment GeoCalib
4.341s 0.577s 1.459s
We evaluate the inference efficiency of Uni3C in comparison to the baseline model, RealisDance-DiT (Zhou et al., 2025b), across two different resolutions, as detailed in Table 6. Note that the efficiency of RealisDance-DiT (Zhou et al., 2025b) is almost the same as the basic VDM, Wan-I2V (Wang et al., 2025a), while only a few convolution layers are newly incorporated to encode additional conditions. For this analysis, we utilize the official settings from Wan2.1 (Wang et al., 2025a), which include a 40-step denoising process and a classifier-free guidance scale of 5. The inference environment is configured using a sequence parallel based on 8 Nvidia H100 GPUs. We further provide other detailed time costs of Uni3C in Table 7. Overall, Uni3C demonstrates impressive efficiency in this setup.
Appendix C Details of Datasets
Table 8. Dataset details of training PCDController. The training datasets include Co3Dv2 (Reizenstein et al., 2021), DL3DV (Ling et al., 2024), RE10K (Zhou et al., 2018), ACID (Liu et al., 2021), Tartainair (Wang et al., 2020), Map-Free-Reloc (Arnold et al., 2022), WildRGBD (Xia et al., 2024), COP3D (Sinha et al., 2023), and UCo3D (Liu et al., 2025). We dynamically sample subsets for each dataset across training epochs.
Indoor Outdoor Object Synthetic Scene Number
Train Validation Epoch
Co3Dv2 $\checkmark$ $\checkmark$ $\checkmark$ 24,437 53 2,252
DL3DV $\checkmark$ $\checkmark$ 9,808 250 19,868
RE10K $\checkmark$ $\checkmark$ 20,259 50 7,528
ACID $\checkmark$ 2,575 20 1,047
Tartainair $\checkmark$ $\checkmark$ $\checkmark$ 2,834 50 2,834
Map-Free-Reloc $\checkmark$ $\checkmark$ 892 18 1,784
WildRGBD $\checkmark$ $\checkmark$ $\checkmark$ 22,105 46 2,300
COP3D $\checkmark$ $\checkmark$ 2,109 20 2,109
UCo3D $\checkmark$ $\checkmark$ $\checkmark$ 161,591 99 10,000
We summarized our training data setting in Table 8. A sampling strategy across epochs ensures balanced sample scales over different data domains. Formally, we pay more attention to the data with high-quality images and complex camera trajectories of real-world scenarios like DL3DV (Ling et al., 2024) and UCo3D (Liu et al., 2025).
Figure 20. Results of Uni3C under the two-person scenario. We operate two different camera trajectories, while their human motions follow the input video.
Appendix D Inference Details of Camera Control
During the inference phase, we begin by extracting the monocular depth of the reference view using Depth-Pro (Bochkovskii et al., 2024). We then establish the foreground depth medium, defined as the rotation radius, to determine the placement of the initial camera. For extracting the foreground mask, we utilize CarveKit  [^2] https://github.com/OPHoperHPO/image-background-remove-tool. In cases where no foreground is detected, the entire image is treated as the foreground. Our camera control is based on two primitives:
  • Rotation. We define the rotation along azimuth and elevation, respectively. The rotation radius is used to control the distance to the foreground, serving as the center of rotation.
  • Translation: We implement translation along the x, y, and z axes. The translation values are constrained within the range of $[0,1.0]$, which are then multiplied by the rotation radius to ensure that translations do not become excessively large or small.
Overall, the camera control mechanism proposed in this paper is flexible and effective enough to handle most downstream tasks.
Appendix E Multiple-Person Discussion
Uni3C builds upon RealisDance-DiT (Zhou et al., 2025b), which is explicitly tailored for single-person scenarios. Notably, we observe that RealisDance-DiT retains functionality even in two-person scenarios, as verified by the case presented in (Zhou et al., 2025b) and visualized in Figure 20. Formally, we separately align two persons’ SMPL-X representations, guided by their respective 3D keypoints. However, injecting additional control signals from more people would hinder the performance. This outcome is unsurprising, as multi-person control is significantly outside the scope of RealisDance-DiT training. As a plug-and-play module, Uni3C can be integrated with other backbones that are natively designed for addressing multi-person control in the future.